OpenStudy (anonymous):

factor 63x^2 - 175 show me how to do it step by step please

OpenStudy (mertsj):

First factor out 7

OpenStudy (mertsj):

Can you do that?

OpenStudy (anonymous):

how where is 7

OpenStudy (mertsj):

It is a factor of 63 and of 175

OpenStudy (mertsj):

7 times 9 = 63 7 times 25 = 175 So factor out the 7

OpenStudy (anonymous):

oh greatest common factor?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

greatest common factor is the highest number that goes into both correct

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

ok i did 7(

OpenStudy (mertsj):

What did you get for the other factor?

OpenStudy (anonymous):

what do you mean i thought its 7

OpenStudy (mertsj):

\[63x^2-175=7(9x^2-25)\]

OpenStudy (mertsj):

Now we see that the other factor is 9x^2-25 which is the difference of two squares. Can you factor 9x^2-25?

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Can you factor x^2-4?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

well 3*3=9

OpenStudy (mertsj):

\[x^2-4=(x-2)(x+2)\]

OpenStudy (mertsj):

\[9x^2-25=(3x-5)(3x+5)\]

OpenStudy (anonymous):

oooh so basically we find the greatest common factor and then find what goes into 9 then put that in the nfind what goes into 25

OpenStudy (mertsj):

Not quite. You factor out the greatest common factor and then you factor the binomial that remains by recognizing that it is the difference of two squares.

OpenStudy (anonymous):

can you help me with this one.. a^4 - a^2 b^2

OpenStudy (mertsj):

Have you studied in your class that \[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

yes i understand what we just went over i get it in my own way now can you help me with the one i just posted b4 u commented just that

OpenStudy (anonymous):

can you help me with this one.. a^4 - a^2 b^2

OpenStudy (mertsj):

first factor out the GCF which is a^2

OpenStudy (anonymous):

wait how

OpenStudy (mertsj):

\[a^4-a^2b^2=a^2(\]

OpenStudy (mertsj):

What goes inside the parenthesis?

OpenStudy (anonymous):

why did u pick the a^2

OpenStudy (mertsj):

Because it is the GCF of a^4 and a^2b^2

OpenStudy (anonymous):

how do you know that that

OpenStudy (mertsj):